OpenStudy (anonymous):

y varies directly with x. When y = 20, x = 50. Find x when y = 36

OpenStudy (lgbasallote):

y varies directly with x so y = kx when y = 20 x = 50 20 = 50k 20/50 = k 2/5 = k find x when y = 36 y =kx 36 = 2/5 x solve for x can you do it?

OpenStudy (anonymous):

90?

OpenStudy (anonymous):

yup.